OpenStudy (hockeychick23):

Probability help please? What is the probability that Annie rolls a sum of 6 and Jose rolls doubles?

OpenStudy (curry):

3/12 * 1/2 = 3/24, i believe.

OpenStudy (curry):

That's only if order of which number comes first doesn't matter though.. :/

OpenStudy (hockeychick23):

when you roll dice isn't there 36 possibilities though?

OpenStudy (curry):

there is, if the order matters.

OpenStudy (curry):

so for instance, if the first number has to be 1, and then so on, then yes, 36 possibilities. if it doesn't matter, you can only have 12 different sums.

OpenStudy (hockeychick23):

yea i think its all 36, so then would it be 1/36* 1/6

OpenStudy (curry):

if it's all 35, then it'd be 1/6*1/6 = 1/36

OpenStudy (curry):

if it's all 36*